OpenStudy (anonymous):

An airplane travels 2100km at a speed of 1000km/h, then encounters a headwind which slows it down to 800km for the next 1300km. What was the average speed of the plane for this trip? (3 marks)

OpenStudy (welshfella):

average speed = total distance travelled / total time taken

OpenStudy (welshfella):

For the first part of the journey he work out the time using t1 = d / v.

OpenStudy (welshfella):

and do the same for the second part first part t1 = 2100/100 hours 2nd part t2 = 1300/800 hours

OpenStudy (welshfella):

so average speed = (2100 + 1300) / (t1 + t2)

OpenStudy (welshfella):

* typo first part = 2100 / 1000 hours